XxDeadxX:

\(\mathbb{\color{gold}{\overset{\bigstar ~~^\bigstar ~~\bigstar}{how do you do this but with hearts}}}\)

janiya45:

@dude

Timmyspu:

@umm @Ultrilliam @dude

KjSaif:

I can't read this my eyes

XxDeadxX:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Timmyspu @umm @Ultrilliam @dude \(\color{#0cbb34}{\text{End of Quote}}\) Why u tagging them I got a question

janiya45:

@Ultrilliam @Ultrilliam @umm

XxDeadxX:

I GOT A QUESTION WHY ARE YOU TAGGING THEM

KjSaif:

whoaa chill.

janiya45:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @XxDeadxX I GOT A QUESTION WHY ARE YOU TAGGING THEM \(\color{#0cbb34}{\text{End of Quote}}\) cas we cant see and

Timmyspu:

@Eiwoh2

janiya45:

iDC HOW MUCH U GET MAD

XxDeadxX:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Timmyspu @Eiwoh2 \(\color{#0cbb34}{\text{End of Quote}}\) Stop spam tagging them

janiya45:

make us

janiya45:

@cpl106

XxDeadxX:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @janiya45 make us \(\color{#0cbb34}{\text{End of Quote}}\) Bet

janiya45:

@sammixboo close this question

Timmyspu:

Im not im just trying to help because i dont know how to help u

cpl106:

i dont know how to do latex sorry

XxDeadxX:

I need help y’all spam tagging the mods they prob helping other ppl so

XxDeadxX:

@supie how do you do the star thing but instead of stars make it haerts

supie:

@Laylalyssa

carmelle:

\(\mathbb{\color{gol}{\overset{♥︎ ~~^\ ♥︎~~\ ♥︎}{I~think~I~did~it}}}\)

janiya45:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @carmelle \(\mathbb{\color{gol}{\overset{♥︎ ~~^\ ♥︎~~\ ♥︎}{I~think~I~did~it}}}\) \(\color{#0cbb34}{\text{End of Quote}}\) pretty

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @janiya45 \(\color{#0cbb34}{\text{Originally Posted by}}\) @carmelle \(\mathbb{\color{gol}{\overset{♥︎ ~~^\ ♥︎~~\ ♥︎}{I~think~I~did~it}}}\) \(\color{#0cbb34}{\text{End of Quote}}\) pretty \(\color{#0cbb34}{\text{End of Quote}}\) thx

XxDeadxX:

Hmm

XxDeadxX:

How u do dat

supie:

Layla showed me this... \(\Huge\bf\color{lightskyblue}{\underset{\heartsuit~~^\bigstar ~~\heartsuit}{\overset{❄️✨💙Supie💙✨❄️}{\bigstar ~~_\heartsuit~~\bigstar}}}\) \(\Huge\bf\color{lightskyblue}{\overset{\heartsuit~~^\bigstar ~~\heartsuit}{\underset{❄️✨💙Supie💙✨❄️}{\bigstar ~~_\heartsuit~~\bigstar}}}\)

XxDeadxX:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie Layla showed me this... \(\Huge\bf\color{lightskyblue}{\underset{\heartsuit~~^\bigstar ~~\heartsuit}{\overset{❄️✨💙Supie💙✨❄️}{\bigstar ~~_\heartsuit~~\bigstar}}}\) \(\Huge\bf\color{lightskyblue}{\overset{\heartsuit~~^\bigstar ~~\heartsuit}{\underset{❄️✨💙Supie💙✨❄️}{\bigstar ~~_\heartsuit~~\bigstar}}}\) \(\color{#0cbb34}{\text{End of Quote}}\) WOAH HOW U DO DAT

XxDeadxX:

\(\Huge\bf\color{lightskyblue}{\underset{\heartsuit~~^\bigstar ~~\heartsuit}{\overset{❄️✨💙cool💙✨❄️}{\bigstar ~~_\heartsuit~~\bigstar}}}\) \(\Huge\bf\color{lightskyblue}{\overset{\heartsuit~~^\bigstar ~~\heartsuit}{\underset{❄️✨💙This~Is~cool💙✨❄️}{\bigstar ~~_\heartsuit~~\bigstar}}}\)

supie:

\(\style{fontFamily:Comic Sans;text-shadow: 1px 1px 2px aqua, 0 0 10px #00f7e6, 0 0 5px blue;color:black;font-size:100px;font-weight:bold;padding:50px;width:500px;height:75%;position:absolute;left:0;z-index:500;text-align:center;} {\Huge\sf\color {} {{Yeah~Nice.}}}\)